# Robustness-Congruent Adversarial Training for Secure Machine Learning Model Updates

### v1

@@ Abstract
Machine-learning models demand for periodic updates to improve their average accuracy, exploiting novel architectures and additional data. However, a newly-updated model may commit mistakes that the previous model did not make. Such misclassifications are referred to as negative flips, and experienced by users as a regression of performance. In this work, we show that this problem also affects robustness to adversarial examples, thereby hindering the development of secure model update practices. In particular, when updating a model to improve its adversarial robustness, some previously-ineffective adversarial examples may become misclassified, causing a regression in the perceived security of the system. We propose a novel technique, named robustness-congruent adversarial training, to address this issue. It amounts to fine-tuning a model with adversarial training, while constraining it to retain higher robustness on the adversarial examples that were correctly classified before the update. We show that our algorithm and, more generally, learning with non-regression constraints, provides a theoretically-grounded framework to train consistent estimators. Our experiments on robust models for computer vision confirm that (i) both accuracy and robustness, even if improved after model update, can be affected by negative flips, and (ii) our robustness-congruent adversarial training can mitigate the problem, outperforming competing baseline methods.
@@ 1 Introduction
Many modern applications of machine learning require frequent model updates to keep pace with the introduction of novel and more powerful architectures, as well as with changes in the underlying data distribution. For instance, when dealing with cybersecurity-related tasks like malware detection, novel threats are discovered at a high pace, and machine learning models need to be constantly retrained to learn to detect them with high accuracy. Another example is given by image tagging, in which image classification and detection models are used to tag pictures of users, and the variety of depicted objects and scenarios varies over time, requiring constant updates. In both cases, as novel and more powerful machine learning architectures emerge, they are rapidly adopted to improve the average system performance; consider, for instance, the need for transitioning from convolutional neural networks to transformer-based architectures.
@@
Within the aforementioned scenarios, the practice of delivering frequent model updates opens up a new challenge related to the maintenance of machine learning models and their performance as perceived by the end users. The issue is that average accuracy is not elaborate enough to also account for sample-wise performance. In particular, even if average accuracy increases after update, some samples that were correctly predicted by the previous model might be misclassified after model update. There is indeed no guarantee that a newly-updated model with higher average accuracy will not commit any mistake on the samples that were correctly predicted by the previous model. The samples that were correctly predicted by the previous model and become misclassified after update have been referred to as Negative Flips (NFs) in [1]. Such mistakes are perceived by end users and practitioners as a regression of performance, similarly to what happens in classical software development, where the term “regression” refers to the deterioration of performance after an update. For this reason, reducing negative flips when performing model updates can be considered as important in practice as improving the overall system accuracy. To better understand the relevance of this issue, consider again the case of malware detection. Experiencing NFs in this domain amounts to having either previously-known legitimate samples misclassified as false positives, or previously-detected malware samples misclassified as legitimate, increasing the likelihood of infecting devices in the wild. [^1] https://www.mandiant.com/resources/blog/churning-out-machine-learning-models-handling-changes-in-model-predictions Similarly, for the case of image tagging, users might find some of their photos changing labels and potentially being mislabeled after a model update, resulting in a bad user experience. Yan et al. [1] have been the first to highlight this issue, and proposed an approach aimed at minimizing negative flips, referred to as Positive Congruent Training (PCT). The underlying idea of their method is to include an additional knowledge-distillation loss while training the updated model to retain the behavior of the old model on samples that were correctly classified before model update. This forces the updated model to reduce the number of errors on such samples (i.e., the number of negative flips), while also aiming to improve the average classification accuracy.
@@
In this work, we argue that model updates may not only induce a perceived regression of classification accuracy via negative flips, but also a regression of other trustworthiness-related metrics, including adversarial robustness. Adversarial robustness is the ability of machine learning models to withstand adversarial examples, i.e., inputs carefully perturbed to mislead classification at test time. Recent progress has shown that adversarial robustness can be improved by adopting more recent neural network architectures and data augmentation techniques [2]. When updating the system with a more recent and robust model, one is expected to gain an overall increase in robust accuracy, i.e., the fraction of adversarial examples (crafted within a given perturbation budget) that are correctly classified by the model. However, similarly to the case of classification accuracy, some previously-ineffective adversarial examples may be able to evade the newly-updated model, thereby causing a perceived regression of robustness. We refer to these newly-induced mistakes as robustness negative flips (RNFs) in the remainder of this manuscript. We discuss the different types of regression in machine learning in Sect. 2.
@@
Our contribution is twofold. We are the first to show that the update of robust machine-learning models can cause a perceived regression of their robustness, meaning that the new model may be fooled by adversarial examples which the previous model was robust to. Furthermore, we propose a novel technique named robustness-congruent adversarial training (RCAT) to overcome this issue, presented in Sect. 3. As the name suggests, our methodology utilizes the well-known adversarial training (AT) procedure to update machine learning models by incorporating adversarial examples within their training data [3]. In particular, we enrich AT by re-formulating the optimization problem with an additional non-regression penalty term that forces the model to retain high robustness on previously-detected adversarial examples, minimizing the fraction of robustness negative flips (as well as the fraction of negative flips). Finally, we show that our technique is also theoretically grounded, demonstrating that learning with a non-regression constraint provides a statistically-consistent estimator, without even affecting the usual convergence rate of $O(\nicefrac{{1}}{{\sqrt{n}}})$, where $n$ is the number of training samples.
@@
Our experiments, reported in Sect. 4, confirm the presence of regression when updating robust models, showing that state-of-the-art image classifiers [2] with higher robustness than their predecessors are misled by some previously-detected adversarial examples. We further show that, when updating models using our RCAT approach, we are able to improve the accuracy and robustness of the updated models while containing both negative flips and robustness negative flips, outperforming competing baselines. We discuss related work on backward compatibility of machine-learning models and continual learning in Sect. 5, and conclude the paper by discussing limitations of the current approach and future research directions in Sect. 6.
@@ 2 Regression of Machine Learning Models
Figure 1: Regression modes in machine-learning model updates. Left: Regression of accuracy induced by negative flips (NFs). When updating an old model $f^{\rm old}$ (dashed black line) with a new model $f^{\rm new}$ (solid black line), a test sample $\boldsymbol{x}$ that was correctly classified by $f^{\rm old}$ may be misclassified by $f^{\rm new}$, causing an NF. Right: Regression of robustness induced by robustness negative flips (RNFs). In a different setting, the test sample $\boldsymbol{x}$ may be still correctly classified by $f^{\rm new}$. However, while no adversarial examples are found against $f^{\rm old}$ (since the perturbation domain $\mathcal{B}(\boldsymbol{x})$, represented by the dashed gray box around $\boldsymbol{x}$, never intersects the decision boundary of $f^{\rm old}$), the adversarial example $\boldsymbol{x}^{\prime}$ is misclassified by $f^{\rm new}$, causing an RNF.
Before delving into the discussion about the types of regression which may be experienced when updating machine learning models, we define some basic notation that will be used throughout this manuscript.
Notation. Let us denote with $\mathcal{D}=(\boldsymbol{x}_{i},y_{i})_{i=1}^{n}$ the training set, consisting of $n$ $d$-dimensional samples $\boldsymbol{x}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, [^2] When data is normalized, as for images, typically $\boldsymbol{x}\in\mathcal{X}=[0,1]^{d}$. along with their class labels $y_{i}\in\mathcal{Y}=\{1,\ldots,c\}$. We similarly define the test set as $\mathcal{T}=(\boldsymbol{x}_{j},y_{j})_{j=1}^{m}$, and assume that both datasets are sampled from the same, unknown probability distribution. Accordingly, a machine learning model can be represented as a function $f:\mathcal{X}\mapsto\mathbb{R}^{c}$, which outputs a confidence value for each class (also referred to as logit). For compactness, we will also use $f_{k}(\boldsymbol{x})$ to denote the confidence value associated to class $k$. The predicted class label can be thus denoted with $\hat{y}(\boldsymbol{x})=\arg\max_{k\in\mathcal{Y}}f_{k}(\boldsymbol{x})$, where the dependency of $\hat{y}$ on $\boldsymbol{x}$ may be omitted in some cases to keep the notation uncluttered. Training a model amounts to finding a function $f$ within a feasible set $\mathcal{F}$ (e.g., constrained by a fixed network architecture) that minimizes a loss $L(y,f(\boldsymbol{x}))$ on the training samples in $\mathcal{D}$, typically using a convex loss that provides an upper bound on the classification error (i.e., the zero-one loss), while penalizing complex solutions via a regularization term that prevents overfitting. This helps to find solutions that achieve lower error on the test set $\mathcal{T}$ and are thus expected to generalize better.
@@ 2.1 Regression of Accuracy: Negative Flips (NFs)
Training a model by minimizing the aforementioned objective means finding a suitable set of parameters for $f$ that reduces the probability of misclassification over all samples. However, this formulation does not constrain the behavior of the model on specific samples, or in specific regions of the feature space. Thus, when updating an old model $f^{\rm old}$ with a new model $f^{\rm new}$ that is trained on the same data $\mathcal{D}$ but using a different network architecture, one may observe that $f^{\rm new}$ yields a lower classification error on $\mathcal{T}$ while misclassifying different samples than those that were misclassified by $f^{\rm old}$. In other words, there might be some test samples $\boldsymbol{x}$ for which $\hat{y}^{\rm old}(\boldsymbol{x})=y$ and $\hat{y}^{\rm new}(\boldsymbol{x})\neq y$, being $y$ their true class. These samples are referred to as negative flips (NFs), as they are responsible for worsening the performance of the new model when compared to the old one. Conversely, the samples that were misclassified by the old model but not by the new one are referred to as positive flips as they improve the overall accuracy of $f^{\rm new}$ on $\mathcal{T}$. More formally, we can define the NF rate (measured on the test set $\mathcal{T}$) as suggested by Yan et al. [1]:
@@ 2.1.1 Positive-Congruent Training (PCT)
To reduce the presence of NFs after model update, and prevent the system users from experiencing a regression of accuracy when using the new model, Yan et al. [1] proposed a technique named Positive-Congruent Training (PCT). The underlying idea of PCT is to add a knowledge-distillation term to the loss function optimized when training the new model $f^{\rm new}$. This term is referred to as focal distillation, as it forces $f^{\rm new}$ to produce similar outputs to those of $f^{\rm old}$ on the samples that were correctly classified by the old model. The PCT approach can be formally described as:
@@
While PCT has empirically proven to reduce NFs, it has not been designed to deal with the regression of robustness. Furthermore, its theoretical properties have not been analyzed, and it remains thus unclear whether it provides a sound, statistically-consistent estimator.
@@ 2.2 Regression of Robustness: Robustness Negative Flips (RNFs)
As shown in Fig. 1 (right), the adversarial example $\boldsymbol{x}^{\prime}$ is optimized within an $\ell_{\infty}$-norm (box) constraint of radius $\varepsilon$ centered on the source sample $\boldsymbol{x}$. It is not difficult to see that, while no adversarial example can be found against the old model $f^{\rm old}$ (as the box constraint never intersects its decision boundary), the same does not hold for the new model $f^{\rm new}$, which is evaded by the adversarial example $\boldsymbol{x}^{\prime}$. Accordingly, when updating the old model $f^{\rm old}$ with $f^{\rm new}$, even if the overall robust accuracy may increase, it may happen that the $f^{\rm new}$ is evaded by some adversarial examples that were not found against $f^{\rm old}$, causing what we call robustness negative flips (RNFs). More formally, we define the RNF rate as:
@@
To conclude, note that an input sample $\boldsymbol{x}$ may be correctly classified by $f^{\rm old}$, and at the same time no corresponding adversarial example may exist, i.e., no $\boldsymbol{x}^{\prime}$ evading $f^{\rm old}$ can be found within the given perturbation domain $\mathcal{B}(\boldsymbol{x})$, as shown in the right plot of Fig. 1. After update, it may happen that $f^{\rm new}$ misclassifies the input sample $\boldsymbol{x}$. Thus, by definition, $\boldsymbol{x}$ coincides with the solution of Problem (5), i.e., its adversarial example $\boldsymbol{x}^{\prime}$. This case will be accounted both as an NF and as an RNF. However, as we will show in our experiments, these joint negative flips are typically very rare and overall negligible.
@@ 3.1 Robustness-Congruent Adversarial Training
This formulation presents two main changes with respect to PCAT (Problem 7), to overcome the two aforementioned limitations of such method. First, to facilitate hyperparameter tuning, we redefine the range of the hyperparameters $\alpha,\beta\in[0,1]$, while fixing $\gamma=1-\alpha-\beta$, so that the three hyperparameters sum up to 1. We also remove the hyperparameter $\lambda$ as it is redundant. Second, we use $f^{\rm src}$ instead of $f^{\rm old}$ in the $\alpha$-scaled term of the focal distillation loss (Eq. 3). The reason is that one may want to update a model $f^{\rm old}$ with an already-trained source model $f^{\rm src}$ that exhibits improved accuracy and robustness, while also reducing NFs and RNFs after update. As $f^{\rm src}$ can be used to initialize $f$ before training via RCAT, it is not reasonable to enforce the behavior of $f^{\rm old}$ over the whole input space. We can indeed try to preserve the behavior of the improved $f^{\rm src}$ model over the whole input space, while enforcing that of $f^{\rm old}$ only on those regions of the input space in which no adversarial examples against $f^{\rm old}$ are found. This is especially convenient when dealing with robust models, as they are normally trained with complicated variants of adversarial training and data augmentation, resulting in a significant computational complexity increase. Thus, if an already-trained, more robust model becomes available, it can be readily used in RCAT as $f^{\rm src}$, as well as to initialize $f$ before optimizing it, while $f^{\rm old}$ can be used to reduce NFs and RNFs in the $\beta$-scaled term.
@@
    it gives the possibility of distilling from a different model than $f^{\rm old}$ over the whole input space, when available, to retain better accuracy and robustness.
@@
$f_{\boldsymbol{w}}\leftarrow f^{\rm src}$   normal-$\triangleright$ Initialize the model as $f^{\rm src}$
1 for $i\in[1,E]$ do
@@
             $\boldsymbol{x}^{\star}\leftarrow\operatornamewithlimits{\arg\,\max}_{\boldsymbol{x}^{\prime}\in\mathcal{B}(\boldsymbol{x})}\mathcal{L}(y,f_{\boldsymbol{w}}(\boldsymbol{x}^{\prime}))$   normal-$\triangleright$ Compute the adversarial example $\boldsymbol{x}^{\star}$ with attack $a$.
             $\boldsymbol{w}\leftarrow\boldsymbol{w}-\eta\nabla_{\boldsymbol{w}}\mathcal{L}(y,f_{\boldsymbol{w}}(\boldsymbol{x}^{\star}))$   normal-$\triangleright$ Update the model parameters $\boldsymbol{w}$.
3
4
return $f^{\rm new}\leftarrow f_{\boldsymbol{w}}$   normal-$\triangleright$ Return the RCAT model.

### v2

@@ Abstract
Machine-learning models demand periodic updates to improve their average accuracy, exploiting novel architectures and additional data. However, a newly updated model may commit mistakes the previous model did not make. Such misclassifications are referred to as negative flips, experienced by users as a regression of performance. In this work, we show that this problem also affects robustness to adversarial examples, hindering the development of secure model update practices. In particular, when updating a model to improve its adversarial robustness, previously ineffective adversarial attacks on some inputs may become successful, causing a regression in the perceived security of the system. We propose a novel technique, named robustness-congruent adversarial training, to address this issue. It amounts to fine-tuning a model with adversarial training, while constraining it to retain higher robustness on the samples for which no adversarial example was found before the update. We show that our algorithm and, more generally, learning with non-regression constraints, provides a theoretically-grounded framework to train consistent estimators. Our experiments on robust models for computer vision confirm that both accuracy and robustness, even if improved after model update, can be affected by negative flips, and our robustness-congruent adversarial training can mitigate the problem, outperforming competing baseline methods.
@@ 1 Introduction
Many modern machine learning applications require frequent model updates to keep pace with the introduction of novel and more powerful architectures, as well as with changes in the underlying data distribution. For instance, when dealing with cybersecurity-related tasks like malware detection, novel threats are discovered at a high pace, and machine learning models need to be constantly retrained to learn to detect them with high accuracy. Another example is given by image tagging, in which image classification and detection models are used to tag pictures of users, and the variety of depicted objects and scenarios varies over time, requiring constant updates. In both cases, as novel and more powerful machine learning architectures emerge, they are rapidly adopted to improve the average system performance; consider, for instance, the need for transitioning from convolutional neural networks to transformer-based architectures.
@@
Within the aforementioned scenarios, the practice of delivering frequent model updates opens up a new challenge related to the maintenance of machine learning models and their performance as perceived by the end users. The issue is that average accuracy is not elaborate enough to also account for sample-wise performance. In particular, even if average accuracy increases after an update, some samples that were correctly predicted by the previous model might be misclassified after the model update. There is indeed no guarantee that a newly updated model with higher average accuracy will not commit any mistake on the samples that the previous model correctly predicted. The samples that the previous model correctly predicted and became misclassified after the update have been referred to as Negative Flips (NFs) in [1]. Such mistakes are perceived by end users and practitioners as a regression of performance, similarly to what happens in classical software development, where the term “regression” refers to the deterioration of performance after an update. For this reason, reducing negative flips when performing model updates can be considered as important in practice as improving the overall system accuracy. To better understand the relevance of this issue, consider again the case of malware detection. Experiencing NFs in this domain amounts to having either previously known legitimate samples misclassified as false positives, or previously-detected malware samples misclassified as legitimate, increasing the likelihood of infecting devices in the wild. [^1] https://www.mandiant.com/resources/blog/churning-out-machine-learning-models-handling-changes-in-model-predictions Similarly, for the case of image tagging, users might find some of their photos changing labels and potentially being mislabeled after a model update, resulting in a bad user experience. Yan et al. [1] have been the first to highlight this issue, and proposed an approach aimed at minimizing negative flips, referred to as Positive Congruent Training (PCT). The underlying idea of their method is to include an additional knowledge-distillation loss while training the updated model to retain the behavior of the old model on samples that were correctly classified before the model update. This forces the updated model to reduce the number of errors on such samples (i.e., the NFs), while also aiming to improve the average classification accuracy.
@@
In this work, we argue that model updates may not only induce a perceived regression of classification accuracy via negative flips, but also a regression of other trustworthiness-related metrics, including adversarial robustness. Adversarial robustness is the ability of machine learning models to withstand adversarial examples, i.e., inputs carefully perturbed to mislead classification at test time [2, 3]. Recent progress has shown that adversarial robustness can be improved by adopting more recent neural network architectures and data augmentation techniques [4]. When updating the system with a more recent and robust model, one is expected to gain an overall increase in robust accuracy, i.e., the fraction of samples for which no adversarial example (crafted within a given perturbation budget) is found. However, similarly to the case of classification accuracy, previously ineffective adversarial attacks on some samples may be able to evade the newly updated model, thereby causing a perceived regression of robustness. In the remainder of this manuscript, we refer to these newly induced mistakes as robustness negative flips (RNFs). We discuss the different types of regression in machine learning in Sect. 2.
@@
Our contribution is twofold. We are the first to show that the update of robust machine-learning models can cause a perceived regression of their robustness, meaning that the new model may be fooled by adversarial examples that the previous model was robust to. Furthermore, we propose a novel technique named robustness-congruent adversarial training (RCAT) to overcome this issue, presented in Sect. 3. As the name suggests, our methodology utilizes the well-known adversarial training (AT) procedure to update machine learning models by incorporating adversarial examples within their training data [5]. In particular, we enrich AT by re-formulating the optimization problem with an additional non-regression penalty term that forces the model to retain high robustness on the training samples for which no adversarial example was found, minimizing the fraction of robustness negative flips (as well as the fraction of negative flips). Finally, we show that our technique is also theoretically grounded, demonstrating that learning with a non-regression constraint provides a statistically-consistent estimator, without even affecting the usual convergence rate of $O(\nicefrac{{1}}{{\sqrt{n}}})$, where $n$ is the number of training samples.
@@
Our experiments, reported in Sect. 4, confirm the presence of regression when updating robust models, showing that state-of-the-art image classifiers [4] with higher robustness than their predecessors are misled by some previously-detected adversarial examples. We further show that, when updating models using RCAT, the accuracy and robustness of the updated models improve while containing both negative flips and robustness negative flips, outperforming competing baselines. We discuss related work on backward compatibility of machine-learning models and continual learning in Sect. 5, and conclude the paper by discussing limitations of the current approach and future research directions in Sect. 6. The code to reproduce all experiments and results is available at: https://github.com/pralab/robustness-congruent-advtrain.
@@ 2 Regression of Machine Learning Models
Figure 1: Regression modes in machine-learning model updates. Left: Regression of accuracy induced by negative flips (NFs). When updating an old model $f^{\rm old}$ (dashed black line) with a new model $f^{\rm new}$ (solid black line), a test sample $\boldsymbol{x}$ that was correctly classified by $f^{\rm old}$ may be misclassified by $f^{\rm new}$, causing an NF. Right: Regression of robustness induced by robustness negative flips (RNFs). In a different setting, the test sample $\boldsymbol{x}$ may still be correctly classified by $f^{\rm new}$. However, while no adversarial examples are found against $f^{\rm old}$ (since the perturbation domain $\mathcal{B}(\boldsymbol{x})$, represented by the dashed gray box around $\boldsymbol{x}$, never intersects the decision boundary of $f^{\rm old}$), an adversarial example $\boldsymbol{x}^{\prime}$ is found against $f^{\rm new}$, causing an RNF.
Before delving into the discussion about the types of regression that may be experienced when updating machine learning models, we define some basic notation that will be used throughout this manuscript.
Notation. Let us denote with $\mathcal{D}=(\boldsymbol{x}_{i},y_{i})_{i=1}^{n}$ the training set, consisting of $n$ $d$-dimensional samples $\boldsymbol{x}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$, [^2] When data is normalized, as for images, typically $\boldsymbol{x}\in\mathcal{X}=[0,1]^{d}$. along with their class labels $y_{i}\in\mathcal{Y}=\{1,\ldots,c\}$. We similarly define the test set as $\mathcal{T}=(\boldsymbol{x}_{j},y_{j})_{j=1}^{m}$, and assume that both datasets are sampled from the same, unknown probability distribution. Accordingly, a machine learning model can be represented as a function $f:\mathcal{X}\mapsto\mathbb{R}^{c}$, which outputs a confidence value for each class (also referred to as logit). For compactness, we will also use $f_{k}(\boldsymbol{x})$ to denote the confidence value associated to class $k$. The predicted class label can be thus denoted with $\hat{y}(\boldsymbol{x})=\arg\max_{k\in\mathcal{Y}}f_{k}(\boldsymbol{x})$, where the dependency of $\hat{y}$ on $\boldsymbol{x}$ may be omitted in some cases to keep the notation uncluttered. Training a model amounts to finding a function $f$ within a feasible set $\mathcal{F}$ (e.g., constrained by a fixed network architecture) that minimizes a loss $L(y,f(\boldsymbol{x}))$ on the training samples in $\mathcal{D}$, typically using a convex loss that provides an upper bound on the classification error (i.e., the zero-one loss), while penalizing complex solutions via a regularization term that prevents overfitting. This helps find solutions that achieve lower error on the test set $\mathcal{T}$ and are thus expected to generalize better.
@@ 2.1 Regression of Accuracy: Negative Flips (NFs)
Training a model by minimizing the aforementioned objective means finding a suitable set of parameters for $f$ that reduces the probability of misclassification over all samples. However, this formulation does not constrain the behavior of the model on specific samples or regions of the feature space. Thus, when updating an old model $f^{\rm old}$ with a new model $f^{\rm new}$ that is trained on the same data $\mathcal{D}$ but using a different network architecture, one may observe that $f^{\rm new}$ yields a lower classification error on $\mathcal{T}$ while misclassifying different samples than those that were misclassified by $f^{\rm old}$. In other words, there might be some test samples $\boldsymbol{x}$ for which $\hat{y}^{\rm old}(\boldsymbol{x})=y$ and $\hat{y}^{\rm new}(\boldsymbol{x})\neq y$, being $y$ their true class. These samples are referred to as negative flips (NFs), as they are responsible for worsening the performance of the new model when compared to the old one. Conversely, the samples that were misclassified by the old model but not by the new one are referred to as positive flips as they improve the overall accuracy of $f^{\rm new}$ on $\mathcal{T}$. More formally, we can define the NF rate (measured on the test set $\mathcal{T}$) as suggested by Yan et al. [1]:
@@ 2.1.1 Positive-Congruent Training (PCT)
To reduce the presence of NFs after model update, and prevent the system users from experiencing a regression of accuracy when using the new model, Yan et al. [1] proposed a technique named Positive-Congruent Training (PCT). The underlying idea of PCT is to add a knowledge-distillation term to the loss function optimized when training the new model $f^{\rm new}$. This term is referred to as focal distillation, as it forces $f^{\rm new}$ to produce outputs similar to those of $f^{\rm old}$ on the samples that the old model correctly classified. The PCT approach can be formally described as:
@@
While PCT has empirically proven to reduce NFs, it has not been designed to deal with the regression of robustness. Furthermore, its theoretical properties have not been analyzed, and it thus remains unclear whether it provides a sound, statistically-consistent estimator.
@@ 2.2 Regression of Robustness: Robustness Negative Flips (RNFs)
As shown in Fig. 1 (right), the adversarial example $\boldsymbol{x}^{\prime}$ is optimized within an $\ell_{\infty}$-norm (box) constraint of radius $\varepsilon$ centered on the source sample $\boldsymbol{x}$. It is not difficult to see that, while no adversarial example can be found against the old model $f^{\rm old}$ (as the box constraint never intersects its decision boundary), the same does not hold for the new model $f^{\rm new}$, which is evaded by the adversarial example $\boldsymbol{x}^{\prime}$. Accordingly, when updating the old model $f^{\rm old}$ with $f^{\rm new}$, even if the overall robust accuracy may increase, $f^{\rm new}$ might be evaded by some adversarial examples that were not found against $f^{\rm old}$, causing what we call robustness negative flips (RNFs). More formally, we define the RNF rate as:
@@
To conclude, note that an input sample $\boldsymbol{x}$ may be correctly classified by $f^{\rm old}$, and at the same time no corresponding adversarial example may exist, i.e., no $\boldsymbol{x}^{\prime}$ evading $f^{\rm old}$ can be found within the given perturbation domain $\mathcal{B}(\boldsymbol{x})$, as shown in the right plot of Fig. 1. After the update, it may happen that $f^{\rm new}$ misclassifies the input sample $\boldsymbol{x}$. Thus, by definition, $\boldsymbol{x}$ coincides with the solution of Problem (5), i.e., its adversarial example $\boldsymbol{x}^{\prime}$. This case will be accounted for as an NF and an RNF simultaneously. However, as discussed in our experiments, these joint negative flips are typically very rare and overall negligible.
@@ 3.1 Robustness-Congruent Adversarial Training
This formulation presents two main changes with respect to PCAT (Problem 7), to overcome the two aforementioned limitations of such method. First, to facilitate hyperparameter tuning, we redefine the range of the hyperparameters $\alpha,\beta\in[0,1]$, while fixing $\gamma=1-\alpha-\beta$, so that the three hyperparameters sum up to 1. We also remove the hyperparameter $\lambda$ as it is redundant. Second, we use $f^{\rm src}$ instead of $f^{\rm old}$ in the $\alpha$-scaled term of the focal distillation loss (Eq. 3). The reason is that one may want to update a model $f^{\rm old}$ with an already-trained source model $f^{\rm src}$ that exhibits improved accuracy and robustness, while also reducing NFs and RNFs after update. As $f^{\rm src}$ can be used to initialize $f$ before training via RCAT, it is not reasonable to enforce the behavior of $f^{\rm old}$ over the whole input space. We can indeed try to preserve the behavior of the improved $f^{\rm src}$ model over the whole input space, while enforcing that of $f^{\rm old}$ only on those regions of the input space in which no adversarial examples against $f^{\rm old}$ are found. This is especially convenient when dealing with robust models, as they are usually trained with complicated adversarial training and data augmentation variants, resulting in a significant increase in computational complexity. Thus, if an already-trained, more robust model becomes available, it can be readily used in RCAT as $f^{\rm src}$, as well as to initialize $f$ before optimizing it, while $f^{\rm old}$ can be used to reduce NFs and RNFs in the $\beta$-scaled term.
@@
    it allows distilling from a different model than $f^{\rm old}$ over the whole input space, when available, to retain better accuracy and robustness.
@@
1
$f_{\boldsymbol{w}}\leftarrow f^{\rm src}$
  $\triangleright$ Initialize the model as $f^{\rm src}$
2
3for $i\in[1,E]$ do
@@
             $\boldsymbol{x}^{\star}\leftarrow\operatornamewithlimits{\arg\,\max}_{\boldsymbol{x}^{\prime}\in\mathcal{B}(\boldsymbol{x})}\mathcal{L}(y,f_{\boldsymbol{w}}(\boldsymbol{x}^{\prime}))$
              $\triangleright$ Compute the adversarial example $\boldsymbol{x}^{\star}$ with attack $a$.
5
            $\boldsymbol{w}\leftarrow\boldsymbol{w}-\eta\nabla_{\boldsymbol{w}}\mathcal{L}(y,f_{\boldsymbol{w}}(\boldsymbol{x}^{\star}))$
              $\triangleright$ Update the model parameters $\boldsymbol{w}$.
6
7
return $f^{\rm new}\leftarrow f_{\boldsymbol{w}}$
  $\triangleright$ Return the RCAT model.